OpenStudy (anonymous):

2/3+n+6=3/4 help me find n

OpenStudy (anonymous):

do 2/3 + 6 = 20/3 so n+20/3 = 3/4 3/4 - 20/3 = -71/12

OpenStudy (anonymous):

i just did it